OpenStudy (anonymous):

the product of 2 positive intergers is 90.find the 2 intergers if the larger is 3 more than twice the smaller show work

OpenStudy (anonymous):

xy=90 Let x be larger x=3+2y x=90/y 3+2y=90/y Solve it

OpenStudy (anonymous):

x*(3+2x)=90